### v2

Abstract.
The aim of this paper is to study $m$-isometric weighted shifts with operator weights (both unilateral and bilateral). We obtain a characterization of such shifts by polynomials with operator coefficients. The procedure of construction of all $m$-isometric weighted shifts with positive weights is presented. We answer the question when the weights of $m$-isometric shift are commuting. The completion problem for $m$-isometric weighted shifts with operator weights is solved. We characterize $m$-isometric bilateral shifts the adjoints of which are also $m$-isometric.
1. Introduction
Classical weighted shifts (both unilateral and bilateral) form an important class of operators on Hilbert space, which is a rich source of examples in operator theory. They have been studied extensively for many years (see [23] for comprehensive survey in this topic). In this paper we deal with a certain generalization of such operators, namely weighted shifts with operator weights. We consider unilateral weighted shifts:
$\ell^{2}(\mathbb{N},H)\ni(h_{0},h_{1},\ldots)\longmapsto(0,S_{1}h_{0},S_{2}h_{1},\ldots)\in\ell^{2}(\mathbb{N},H)$
and bilateral weighted shifts:
$\ell^{2}(\mathbb{Z},H)\ni(\ldots,h_{-1},\boxed{h_{0}},h_{1},\ldots)\longmapsto(\ldots,S_{-1}h_{-2},\boxed{S_{0}h_{-1}},S_{1}h_{0},\ldots)\in\ell^{2}(\mathbb{Z},H),$
[…]
$m$-isometric operators were introduced by Agler in [2] and investigated further in [3] (and its subsequent parts). From that time on $m$-isometries are subject of long-term research in operator theory (see [9], [14], [15], [24],[6], [7]). In [13] Jabłoński characterized 2-isometric unilateral weighted shifts with operator weights. In [4], on the occasion of solving the Cauchy dual subnormality problem for 2-isometries, Anand et.al. showed that a non-unitary 2-isometric operator satisfying so-called kernel condition is (up to a unitary equivalence) the orthogonal sum of unitary operator and unilateral weighted shift with positive and invertible weights (see [16] for certain generalization of this fact in case of general $m$-isometries).
$m$-isometric weighted shifts (with scalar weights) were characterized first by Bermúdez et.al. in [8]. Some time later, in [1] there appeared a more convenient characterization of $m$-isometric shifts by real polynomials of degree at most $m-1$. Our goal is to generalize the results of [1]. The paper is organized as follows. In Section 2 we set up notation and terminology and review some of the standard facts on weighted shifts with operator weights. Section 3 is devoted to study $m$-isometric unilateral weighted shifts. Theorem 3.1 characterizes $m$-isometric unilateral weighted shifts with operator weights by polynomials of degree at most $m-1$ with operator coefficients; it may be viewed as a certain generalization of [1, Theorem 1]. In Theorem 3.4 we give the method of constructing all $m$-isometric weighted shifts with positive and invertible weights. Theorem 3.10 answers the question when the weights of $m$-isometric unilateral weighted shifts commute; it turns out that they commute if and only if the coefficients of polynomial given by Theorem 3.1 commute. We also solve the completion problem for $m$-isometric unilateral weighted shifts with operator weights (see Theorem 3.9). Several examples are also provided. The aim of Section 4 is to provide counterparts of results presented in Section 3 for bilateral weighted shifts. Beyond that it is shown that the adjoint of $m$-isometric bilateral weighted shift with operator weights is also $m$-isometric if and only if the polynomial given by Theorem 4.4 is invertible (as an element of the algebra of polynomials with operator coefficients). This allows us to answer the question in [7, Problem 4.9] in negative.
2. Preliminaries
By $\mathbb{Z}$ we denote the set of integers and by $\mathbb{N}$ the set of non-negative integers. For $p\in\mathbb{Z}$ we define
$\mathbb{Z}_{p}=\left\{n\in\mathbb{Z}\colon n\geqslant p\right\}.$
We denote by $\mathbb{R}$ and $\mathbb{C}$ the fields of real and complex numbers, respectively. For simplicity of notation we write $(a_{i})_{i\in I}\subset A$, if $(a_{i})_{i\in I}$ is the sequence of elements of a set $A$. If $A$ is $n\times n$ matrix, then $\mathrm{tr}\,M$ stands for the trace of $M$.
If $R$ is an algebra over a field $\mathbb{K}\in\left\{\mathbb{R},\mathbb{C}\right\}$, then $R[z]$ stands for the algebra of all polynomials with coefficients in $R$ with standard addition and multiplication. For $n\in\mathbb{N}$, $R_{n}[z]$ denotes the vector space over $\mathbb{K}$ of all polynomials with coefficients in $R$ of degree less than or equal to $n$.
Throughout this paper all Hilbert spaces are assumed to be complex. Let $H$ be a Hilbert space. By $\mathbf{B}(H)$ we denote the $C^{\ast}$-algebra of all linear and bounded operators on $H$. For $T\in\mathbf{B}(H)$, we will write $\mathcal{N}(T)$ for the kernel and $\mathcal{R}(T)$ for the range of $T$; by $\sigma(T)$ and $r(T)$ we denote the spectrum and the spectral radius of $T$, respectively. For $T\in\mathbf{B}(H)$ we have the following so-called range-kernel decomposition:
(2.1) $H=\mathcal{N}(T)\oplus\overline{\mathcal{R}(T^{\ast})}.$
An operator $U\in\mathbf{B}(\mathcal{H})$ is called a partial isometry if $U|_{\mathcal{N}(U)^{\perp}}$ is isometric. For every $T\in\mathbf{B}(\mathcal{H})$ there exists the unique partial isometry $U\in\mathbf{B}(\mathcal{H})$ such that $\mathcal{N}(T)=\mathcal{N}(U)$ and $T=U\lvert T\rvert$, where $\lvert T\rvert=(T^{\ast}T)^{1/2}$ (see [25, Theorem 7.20]); the decomposition $T=U\lvert T\rvert$ is called the polar decomposition of $T$. The operators $T,S\in\mathbf{B}(H)$ are unitarily equivalent if there exists a unitary operator $U\in\mathbf{B}(H)$ such that $US=TU$. Note that if $T$ and $S$ are unitarily equivalent, then so are $T^{\ast}$ and $S^{\ast}$ (by the same unitary operator). The operators $T,S\in\mathbf{B}(H)$ are doubly commuting if $TS=ST$ and $T^{\ast}S=ST^{\ast}$. An operator $T\in\mathbf{B}(H)$ is called $m$-isometry ($m\in\mathbb{Z}_{1}$) if
(2.2) $\beta_{m}(T):=\sum_{k=0}^{m}\binom{m}{k}(-1)^{m-k}T^{\ast k}T^{k}=0;$
$T$ is called positive if $\langle Tf,f\rangle>0$ for every $f\in H$, $f\not=0$. For $N\in\left\{\mathbb{N},\mathbb{Z}\right\}$ we define
$\ell^{2}(N,H)=\{(h_{i})_{i\in N}\subset H\colon\sum_{i\in N}\lVert h_{i}\rVert^{2}<\infty\};$
this is a Hilbert space with the inner product given by the formula
$\langle h,h^{\prime}\rangle=\sum_{i\in N}\langle h_{i},h_{i}^{\prime}\rangle,\quad h=(h_{i})_{i\in N},h^{\prime}=(h_{i}^{\prime})_{i\in N}\in\ell^{2}(N,H).$
Every operator $T\in\mathbf{B}(\ell^{2}(N,H))$ has a matrix representation $[T_{i,j}]_{i,j\in N}$, where $T_{i,j}\in\mathbf{B}(H)$, satisfying the following formula (see [11, Chapter 8]):
$T(x_{i})_{i\in N}=\left(\sum_{j\in N}T_{i,j}x_{j}\right)_{i\in N},\quad(x_{i})_{i\in N}\in\ell^{2}(N,H).$
Observe that if $T=[T_{i,j}]_{i,j\in N},S=[S_{i,j}]_{i,j\in N}\in\mathbf{B}(\ell^{2}(N,H))$, then
(2.3) $T^{\ast}=[T^{\ast}_{j,i}]_{i,j\in N}$
and
(2.4) $TS=\left[\sum_{k\in N}T_{i,k}S_{k,j}\right]_{i,j\in N}.$
If $(S_{i})_{i\in\mathbb{Z}_{1}}\subset\mathbf{B}(H)$ is a uniformly bounded sequence of operators, then we define the unilateral weighted shift $S\in\mathbf{B}(\ell^{2}(\mathbb{N},H))$ with weights $(S_{i})_{i\in\mathbb{Z}_{1}}$ as follows
$S(h_{i})_{i\in\mathbb{N}}=(0,S_{1}h_{0},S_{2}h_{1},\ldots),\quad(h_{i})_{i\in\mathbb{N}}\in\ell^{2}(\mathbb{N},H).$
It is a straightforward computation to prove that (cf. [17, Lemma 2.2])
(2.5) $\lVert S\rVert=\sup_{n\in\mathbb{Z}_{1}}\lVert S_{n}\rVert.$
The matrix representation of $S$ takes the form
$S=\begin{bmatrix}0&0&0&0&\ddots\\ S_{1}&0&0&0&\ddots\\ 0&S_{2}&0&0&\ddots\\ 0&0&S_{3}&0&\ddots\\ \ddots&\ddots&\ddots&\ddots&\ddots\end{bmatrix}.$
Similarly, if $(S_{i})_{i\in\mathbb{Z}}\subset\mathbf{B}(H)$ is a uniformly bounded sequence of operators, we define the bilateral weighted shift $S\in\mathbf{B}(\ell^{2}(\mathbb{Z},H))$ with weights $(S_{i})_{i\in\mathbb{Z}}$ by the formula
$S(h_{i})_{i\in\mathbb{Z}}=(S_{i}h_{i-1})_{i\in\mathbb{Z}},\quad(h_{i})_{i\in\mathbb{Z}}\in\ell^{2}(\mathbb{Z},H);$
the matrix representation of this operator takes the form
(2.6) $S=\begin{bmatrix}\ddots&\ddots&\ddots&\ddots&\ddots\\ \ddots&0&0&0&\ddots\\ \ddots&S_{0}&\boxed{0}&0&\ddots\\ \ddots&0&S_{1}&0&\ddots\\ \ddots&\ddots&\ddots&\ddots&\ddots\end{bmatrix},$
[…]
(2.7) $\lVert S\rVert=\sup_{n\in\mathbb{Z}}\lVert S_{n}\rVert.$
It turns out that the adjoint of a bilateral weighted shift is also (up to a unitary equivalence) a bilateral weighted shift.
Lemma 2.1.
Let $H$ be a Hilbert space. If $S\in\mathbf{B}(\ell^{2}(\mathbb{Z}),H)$ is a bilateral weighted shift with weights $(S_{j})_{j\in\mathbb{Z}}\subset\mathbf{B}(H)$, then $S^{\ast}$ is unitarily equivalent to a bilateral weighted shift with weights $(S_{-j+1}^{\ast})_{j\in\mathbb{Z}}$.
Proof.
By (2.3) and (2.6), we have
$S^{\ast}=\begin{bmatrix}\ddots&\ddots&\ddots&\ddots&\ddots\\ \ddots&0&S_{0}^{\ast}&0&\ddots\\ \ddots&0&\boxed{0}&S_{1}^{\ast}&\ddots\\ \ddots&0&0&0&\ddots\\ \ddots&\ddots&\ddots&\ddots&\ddots\end{bmatrix},$
so $S^{\ast}(x_{j})_{j\in\mathbb{Z}}=(S_{j+1}^{\ast}x_{j+1})_{j\in\mathbb{Z}}$ for $(x_{j})_{j\in\mathbb{Z}}\in\ell^{2}(\mathbb{Z},H)$. Denote by $\widehat{S}$ the bilateral weighted shift with weights $(S_{-j+1}^{\ast})_{j\in\mathbb{Z}}$. Define $U\in\mathbf{B}(\ell^{2}(\mathbb{Z},H))$ by the formula:
$U(x_{j})_{j\in\mathbb{Z}}=(x_{-j})_{j\in\mathbb{Z}},\quad(x_{j})_{j\in\mathbb{Z}}\in\ell^{2}(\mathbb{Z},H).$
Clearly, $U$ is a unitary operator. Moreover,
$\widehat{S}U(x_{j})_{j\in\mathbb{Z}}=\widehat{S}(x_{-j})_{j\in\mathbb{Z}}=(S_{-j+1}^{\ast}x_{-j+1})_{j\in\mathbb{Z}},\quad(x_{j})_{j\in\mathbb{Z}}\in\ell^{2}(\mathbb{Z},H),$
and
$US^{\ast}(x_{j})_{j\in\mathbb{Z}}=U(S_{j+1}^{\ast}x_{j+1})_{j\in\mathbb{Z}}=(S_{-j+1}^{\ast}x_{-j+1})_{j\in\mathbb{Z}},\quad(x_{j})_{j\in\mathbb{Z}}\in\ell^{2}(\mathbb{Z},H),$
so $\widehat{S}U=US^{\ast}$. Therefore, $U$ makes $S^{\ast}$ and $\widehat{S}$ unitarily equivalent. ∎
It can be easily seen that
(2.8) $S\text{ is invertible}\iff S_{j}\text{ is invertible for every }j\in\mathbb{Z}.$
The proof of the following observation is a straightforward application of (2.3) and (2.4) (see also [10, Lemma 2.1])
Observation 2.2.
Let $H$ be a Hilbert space.
  1. (i)
    If $S\in\mathbf{B}(\ell^{2}(\mathbb{N}),H)$ is a unilateral weighted shift with weights $(S_{i})_{i\in\mathbb{Z}_{1}}\subset\mathbf{B}(H)$, then for every $n\in\mathbb{Z}_{1}$,
    (2.9) $(S^{\ast n}S^{n})_{i,j}=\begin{cases}\lvert S_{i+n}\cdots S_{i+1}\rvert^{2},&\text{if }i=j\\ 0,&\text{otherwise}\end{cases},\quad i,j\in\mathbb{N},$
    are the entries of the matrix representation of $S^{\ast n}S^{n}$.
  2. (ii)
    If $S\in\mathbf{B}(\ell^{2}(\mathbb{Z}),H)$ is a bilateral weighted shift with weights $(S_{i})_{i\in\mathbb{Z}}\subset\mathbf{B}(H)$, then for every $n\in\mathbb{Z}_{1}$ the entries of the matrix representation of $S^{\ast n}S^{n}$ are given by (2.9) for all $i,j\in\mathbb{Z}$.
3. $m$-isometric unilateral weighted shifts
In this section we are going to characterize $m$-isometric unilateral weighted shifts with invertible weights. Let us state the main result of this section. It is a natural generalization of [1, Theorem 1].
Theorem 3.1.
Let $H$ be a Hilbert space. Suppose $S\in\mathbf{B}(\ell^{2}(\mathbb{N}),H)$ is a unilateral weighted shift with invertible weights $(S_{j})_{j\in\mathbb{Z}_{1}}$. For $m\in\mathbb{Z}_{1}$ the following conditions are equivalent:
  1. (i)
    $S$ is $m$-isometric,
  2. (ii)
    for every $x\in H$, $x\not=0$, the classical weighted shift $S_{x}\in\ell^{2}(\mathbb{N},\mathbb{C})$ with the weights $(\alpha_{j}(x))_{j\in\mathbb{Z}_{1}}\subset(0,\infty)$ given by
    $\alpha_{j}(x)=\begin{cases}\lVert S_{1}x\rVert,&j=1\\ \frac{\lVert S_{j}\cdots S_{1}x\rVert}{\lVert S_{j-1}\cdots S_{1}x\rVert},&j\geqslant 2\end{cases},$
    is $m$-isometric,
  3. (iii)
    there exists a polynomial $p\in\mathbf{B}(H)_{m-1}[z]$ such that $p(0)=I$ and
    (3.1) $p(n)=\lvert S_{n}\cdots S_{1}\rvert^{2},\quad n\in\mathbb{Z}_{1}.$
Moreover, if (iii) holds, then
(3.2) $\lVert S\rVert^{2}=\sup\left\{\frac{\langle p(n+1)h,h\rangle}{\langle p(n)h,h\rangle}\!:n\in\mathbb{N},h\in H,h\not=0\right\}.$
Before we prove this theorem, we need a technical lemma.
Lemma 3.2.
Let $H$ and $S$ be as in Theorem 3.1. If $p\in\mathbf{B}(H)_{m-1}[z]$ is a polynomial as in Theorem 3.1(iii), then every coefficient of $p$ is a linear combination of operators $I,\lvert S_{1}\rvert^{2},\ldots,\lvert S_{m-1}\cdots S_{1}\rvert^{2}$ with real coefficients. In particular, every coefficient of $p$ is self-adjoint and for every $x\in H$, $\langle p(\,\cdot\,)x,x\rangle\in\mathbb{R}_{m-1}[z]$.
Proof.
Write $p(z)=A_{m-1}z^{m-1}+\ldots+A_{1}z+I$, where $A_{m-1},\ldots,A_{1}\in\mathbf{B}(H)$. Then for every $x\in H$,
$q_{x}(z):=\langle p(z)x,x\rangle=\sum_{j=0}^{m-1}\langle A_{j}x,x\rangle z^{j}\in\mathbb{C}_{m-1}[z].$
By the assumption, the coefficients of $q_{x}$ satisfy the following system of linear equations [^1] We stick to the convention that $10^{0}=$ 1:
$\begin{bmatrix}0^{m-1}&\ldots&0^{0}\\ 1^{m-1}&\ldots&1^{0}\\ \vdots&\vdots&\vdots\\ (m-1)^{m-1}&\ldots&(m-1)^{0}\end{bmatrix}\begin{bmatrix}\langle A_{m-1}x,x\rangle\\ \langle A_{m-2}x,x\rangle\\ \vdots\\ \langle x,x\rangle\end{bmatrix}=\begin{bmatrix}\langle x,x\rangle\\ \langle\lvert S_{1}\rvert^{2}x,x\rangle\\ \vdots\\ \langle\lvert S_{m-1}\cdots S_{1}\rvert^{2}x,x\rangle\end{bmatrix}.$
By the Cramer rule (see [19, Theorem 3.36.(7)]), the only solution of the above system is of the form
$\langle A_{j}x,x\rangle=M_{j,0}\langle x,x\rangle+\sum_{\ell=1}^{m-1}M_{j,\ell}\langle\lvert S_{\ell}\cdots S_{1}\rvert^{2}x,x\rangle,\quad j=0,1,\ldots,m-1,$
where the constants $M_{j,\ell}\in\mathbb{R}$ depend only on $j$ and $\ell$ (not on $x\in H$). Thus,
$A_{j}=M_{j,0}I+\sum_{\ell=1}^{m-1}M_{j,\ell}\lvert S_{\ell}\cdots S_{1}\rvert^{2},\quad j=0,\ldots,m-1.$
This completes the proof. ∎
Proof of Theorem 3.1.
(i)$\iff$(ii). We infer from Observation 2.2 that the matrix of $\beta_{m}(S)$ is diagonal and that the entries on the diagonal of this matrix are given by the formula
$(\beta_{m}(S))_{j,j}=\sum_{k=0}^{m}(-1)^{m-k}\binom{m}{k}\lvert S_{j+k}\cdots S_{j+1}\rvert^{2},\quad j\in\mathbb{N}.$
Therefore, $\beta_{m}(S)=0$ if and only if $(\beta_{m}(S))_{j,j}=0$ for every $j\in\mathbb{N}$. In turn, for $x\in H$, $x\not=0$,
$\langle(\beta_{m}(S))_{j,j}x,x\rangle$ $=\sum_{k=0}^{m}(-1)^{m-k}\binom{m}{k}\lVert S_{j+k}\cdots S_{j+1}x\rVert^{2}$
$=\lVert S_{j}\cdots S_{1}x\rVert^{2}\sum_{k=0}^{m}(-1)^{m-k}\binom{m}{k}\alpha_{j+k}(x)^{2}\cdots\alpha_{0}(x)^{2},$
where $\alpha_{0}(x)=1$ for all $x\in H$. In view of [8, Proposition 3.2], $(\beta_{m}(S))_{j,j}=0$ for every $j\in\mathbb{N}$ if and only if $S_{x}$ is $m$-isometric for every $x\not=0$.
(i)$\Longrightarrow$(iii). By [3, Eq.(1.3)], for every $k\in\mathbb{N}$ we have
(3.3) $S^{\ast k}S^{k}=\sum_{n=0}^{m-1}\frac{1}{n!}\beta_{n}(S)k(k-1)\cdots(k-n+1).$
It follows from Observation 2.2 that
$\lvert S_{k}\cdots S_{1}\rvert^{2}=(S^{\ast k}S^{k})_{0,0}=\sum_{n=0}^{m-1}\frac{1}{n!}\beta_{n}(S)_{0,0}k(k-1)\cdots(k-n+1).$
Writing $k(k-1)\cdots(k-n+1)=\sum_{\ell=0}^{n}C_{n,\ell}k^{\ell}$ ($n\in\mathbb{N}$) we derive from the above equality that
$\lvert S_{k}\cdots S_{1}\rvert^{2}=\sum_{n=0}^{m-1}\frac{1}{n!}\beta_{n}(S)_{0,0}\sum_{\ell=0}^{n}C_{n,\ell}k^{\ell},\quad k\in\mathbb{N}.$
Hence, the polynomial
$p(z)=\sum_{n=0}^{m-1}\frac{1}{n!}\beta_{n}(S)_{0,0}\sum_{\ell=0}^{n}C_{n,\ell}z^{\ell}\in\mathbf{B}(H)_{m-1}[z]$
satisfies (iii).
(iii)$\Longrightarrow$(ii). From (iii) we deduce that for every $x\in H$ and every $n\in\mathbb{Z}_{1}$,
(3.4) $\lVert S_{n}\cdots S_{1}x\rVert^{2}=\langle\lvert S_{n}\cdots S_{1}\rvert^{2}x,x\rangle=\langle p(n)x,x\rangle.$
For $x\in H$ set $q_{x}(z)=\langle p(z)x,x\rangle$. Then, by Lemma 3.2, $q\in\mathbb{R}_{m-1}[x]$ and for $x\not=0$ we have $q(1)=\lVert S_{1}x\rVert^{2}$ and
(3.5) $\frac{\lVert S_{n+1}\cdots S_{1}x\rVert^{2}}{\lVert S_{n}\cdots S_{1}x\rVert^{2}}=\frac{q(n+1)}{q(n)},\quad n\in\mathbb{Z}_{1}.$
Therefore, by [1, Theorem 1], the classical weighted shift $S_{x}$ is $m$-isometric for every $x\not=0$.
Now let us prove the ”moreover” part. Since for every $n\in\mathbb{Z}_{1}$, $S_{n}\cdots S_{1}$ is invertible, we have $\mathcal{R}(S_{n}\cdots S_{1})=H$ and
$\lVert S_{n+1}\rVert^{2}$ =$\sup\left\{\frac{\lVert S_{n+1}h\rVert^{2}}{\lVert h\rVert^{2}}\colon h\in H,h\not=0\right\}$
=$\sup\left\{\frac{\lVert S_{n+1}\cdots S_{1}h\rVert^{2}}{\lVert S_{n}\cdots S_{1}h\rVert^{2}}\colon h\in H,h\not=0\right\}$
=$\sup\left\{\frac{\lVert\lvert S_{n+1}\cdots S_{1}\rvert h\rVert^{2}}{\lVert\lvert S_{n}\cdots S_{1}\rvert h\rVert^{2}}\colon h\in H,h\not=0\right\}$
=$\sup\left\{\frac{\langle\lvert S_{n+1}\cdots S_{1}\rvert^{2}h,h\rangle}{\langle\lvert S_{n}\cdots S_{1}\rvert^{2}h,h\rangle}\colon h\in H,h\not=0\right\}$
=$\sup\left\{\frac{\langle p(n+1)h,h\rangle}{\langle p(n)h,h\rangle}\colon h\in H,h\not=0\right\}.$
Using the fact that $p(0)=I$ and proceeding similarly as above, we also have
$\lVert S_{1}\rVert^{2}=\sup\left\{\frac{\langle p(1)h,h\rangle}{\langle p(0)h,h\rangle}\colon h\in H,h\not=0\right\}.$
Therefore, (3.2) follows from the above and (2.5). ∎
In [1, Theorem 1] the authors actually showed that for a classical unilateral weighted shifts $S\in\mathbf{B}(\ell^{2}(\mathbb{N},\mathbb{C}))$ with scalar weights $(\lambda_{n})_{n\in\mathbb{Z}_{1}}\subset\mathbb{C}\setminus\{0\}$, $m$-isometricity of $S$ is equivalent to the existence of a polynomial $p\in\mathbb{R}_{m-1}[x]$ satisfying the following formula:
$\lvert\lambda_{n}\rvert^{2}=\frac{p(n+1)}{p(n)},\quad n\in\mathbb{Z}_{1}.$
As a corollary from Theorem 3.1, we obtain the counterpart of the above formula for shifts with operator weights.
Corollary 3.3.
Let $H$ and $S$ be as in Theorem 3.1.
  1. (i)
    If $S$ is $m$-isometric and $p\in\mathbf{B}(H)_{m-1}[z]$ is a polynomial as in Theorem 3.1(iii), then for every $n\in\mathbb{Z}_{1}$, $\lvert S_{n}\rvert^{2}$ is unitarily equivalent to $p(n-1)^{-1/2}p(n)p(n-1)^{-1/2}$.
  2. (ii)
    Assume that the weights of $S$ are pairwise doubly commuting. Then $S$ is $m$-isometric if and only if there exists a polynomial $p\in\mathbf{B}(H)_{m-1}[z]$ such that $p(0)=I$ and
    (3.6) $\lvert S_{n}\rvert^{2}=p(n)p(n-1)^{-1}=p(n-1)^{-1}p(n),\quad n\in\mathbb{Z}_{1}.$
Proof.
(i). For $n\in\mathbb{Z}_{1}$ let $S_{n}\cdots S_{1}=U_{n}\lvert S_{n}\cdots S_{1}\rvert$ be the polar decomposition of $S_{n}\cdots S_{1}$. Since $S_{n}\cdots S_{1}$ is invertible, $U_{n}$ is unitary for every $n\in\mathbb{Z}_{1}$. For $n\in\mathbb{Z}_{2}$ we have
$p(n)=\lvert S_{n}\cdots S_{1}\rvert^{2}$ $=(S_{n-1}\cdots S_{1})^{\ast}S_{n}^{\ast}S_{n}S_{n-1}\cdots S_{1}$
$=\lvert S_{n-1}\cdots S_{1}\rvert U_{n-1}^{\ast}\lvert S_{n}\rvert^{2}U_{n-1}\lvert S_{n-1}\cdots S_{1}\rvert$
$=p(n-1)^{1/2}U_{n-1}^{\ast}\lvert S_{n}\rvert^{2}U_{n-1}p(n-1)^{1/2}.$
This implies that
$U_{n-1}^{\ast}\lvert S_{n}\rvert^{2}=p(n-1)^{-1/2}p(n)p(n-1)^{-1/2}U_{n-1}^{\ast},\quad n\in\mathbb{Z}_{2}.$
Since $p(0)=I$, we have
$\lvert S_{1}\rvert^{2}=p(1)=p(0)^{-1/2}p(1)p(0)^{-1/2}.$
Therefore, $\lvert S_{n}\rvert^{2}$ is unitarily equivalent to $p(n-1)^{-1/2}p(n)p(n-1)^{-1/2}$ for every $n\in\mathbb{Z}_{1}$.
(ii). Since the weights of $S$ are doubly commuting, we infer from [5, Lemma 2.1] that
$\lvert S_{n}\cdots S_{1}\rvert^{2}=\lvert S_{n}\rvert^{2}\cdots\lvert S_{1}\rvert^{2},\quad n\in\mathbb{Z}_{1}.$
Using the above equality, we can easily show that a polynomial $p\in\mathbf{B}(H)_{m-1}[z]$ satisfies (3.1) if and only if it satisfies (3.6). ∎
Theorem 3.1 enables us to give a procedure of constructing $m$-isometric unilateral weighted shifts with positive and invertible weights; since, by [17, Theorem 3.4], every unilateral weighted shift with invertible weights is unitarily equivalent to the shift with positive and invertible weights, the assumption of positivity of weights is not restricting.
Theorem 3.4.
Let $H$ be a Hilbert space. If $p\in\mathbf{B}(H)_{m-1}[z]$ is a polynomial such that
  1. (i)
    $p(0)=I$,
  2. (ii)
    $p(n)$ is positive and invertible for every $n\in\mathbb{N}$,
  3. (iii)
    $C(p):=\sup\left\{\frac{\langle p(n+1)h,h\rangle}{\langle p(n)h,h\rangle}\colon n\in\mathbb{N},h\in H,h\not=0\right\}<\infty$,
then there exists an $m$-isometric unilateral weighted shift $S\in\mathbf{B}(\ell^{2}(\mathbb{N},H))$ with positive and invertible weights $(S_{j})_{j\in\mathbb{Z}_{1}}\subset\mathbf{B}(H)$ satisfying (3.1). Moreover, $\lVert S\rVert=\sqrt{C(p)}$.
Proof.
We define the sequence $(S_{j})_{j\in\mathbb{Z}_{1}}$ by induction. Set $S_{1}:=p(1)^{1/2}$. Then $S_{1}$ is positive and invertible and such that $\langle\lvert S_{1}\rvert^{2}h,h\rangle=\langle p(1)h,h\rangle$ for every $h\in H$. Assume now that for some $k\in\mathbb{Z}_{1}$ we have defined the weights $S_{1},\ldots,S_{k}\in\mathbf{B}(H)$ in a way that they are invertible and positive operators satisfying (3.1) for $n=1,\ldots,k$. Then
$\widetilde{S}_{k+1}:=S_{1}^{-1}\cdots S_{k}^{-1}p(k+1)S_{k}^{-1}\cdots S_{1}^{-1}$
is a positive and invertible operator and so is $S_{k+1}:=\widetilde{S}_{k+1}^{1/2}$. We have
$\langle\lvert S_{k+1}\cdots S_{1}\rvert^{2}h,h\rangle=\langle S_{1}\cdots S_{k}\cdot\widetilde{S}_{k+1}\cdot S_{k}\cdots S_{1}h,h\rangle=\langle p(k+1)h,h\rangle,\quad h\in H,$
so (3.1) holds for $n=k+1$. It remains to show that the sequence $(S_{j})_{j\in\mathbb{Z}_{1}}\subset\mathbf{B}(H)$ defined above is uniformly bounded. By (i) and (iii) we have
$\lVert S_{1}h\rVert^{2}=\langle S_{1}^{2}h,h\rangle=\langle p(1)h,h\rangle\leqslant C(p)\cdot\langle p(0)h,h\rangle=C(p)\cdot\lVert h\rVert^{2},\quad h\in H,$
and
$\lVert S_{n+1}\cdots S_{1}h\rVert^{2}$ $=\langle\lvert S_{n+1}\cdots S_{1}\rvert^{2}h,h\rangle=\langle p(n+1)h,h\rangle$
$\leqslant C(p)\cdot\langle p(n)h,h\rangle=C(p)\cdot\lVert S_{n}\cdots S_{1}h\rVert^{2},\quad n\in\mathbb{Z}_{1},h\in H.$
Since $\mathcal{R}(S_{n}\cdots S_{1})=H$ for every $n\in\mathbb{Z}_{1}$, it follows that $\lVert S_{n}\rVert^{2}\leqslant C(p)$ for every $n\in\mathbb{Z}_{1}$. We infer from Theorem 3.1 that the weighted shift $S\in\mathbf{B}(\ell^{2}(\mathbb{N},H))$ with weights $(S_{j})_{j\in\mathbb{Z}_{1}}$ is $m$-isometric. The formula $\lVert S\rVert=\sqrt{C(p)}$ follows from (3.2). ∎
The following simple observation gives us another way of checking the condition Theorem 3.4(iii).
Observation 3.5.
Let $H$ be a Hilbert space. If $p\in\mathbf{B}(H)_{m-1}[z]$ is a polynomial satisfying Theorem 3.4(i)–(ii), then
(3.7) $C(p)=\sup\left\{r(p(n+1)p(n)^{-1})\colon n\in\mathbb{N}\right\}.$
Proof.
By the fact that for every $n\in\mathbb{N}$, $p(n)$ is positive and invertible, we have
$\sup\left\{\frac{\langle p(n+1)h,h\rangle}{\langle p(n)h,h\rangle}\colon n\in\mathbb{N},h\in H,h\not=0\right\}$
=$\sup\left\{\frac{\langle p(n+1)p(n)^{-1/2}h,p(n)^{-1/2}h\rangle}{\langle p(n)p(n)^{-1/2}h,p(n)^{-1/2}h\rangle}\colon n\in\mathbb{N},h\in H,h\not=0\right\}$
=$\sup\left\{\frac{\langle p(n)^{-1/2}p(n+1)p(n)^{-1/2}h,h\rangle}{\langle h,h\rangle}\colon n\in\mathbb{N},h\in H,h\not=0\right\}$
$=\sup\left\{\lVert p(n)^{-1/2}p(n+1)p(n)^{-1/2}\rVert\colon n\in\mathbb{N}\right\}.$
Since $p(n)^{-1/2}p(n+1)p(n)^{-1/2}$ is positive for every $n\in\mathbb{N}$, it follows from [25, Theorem 5.44] that
$\lVert p(n)^{-1/2}p(n+1)p(n)^{-1/2}\rVert=r(p(n)^{-1/2}p(n+1)p(n)^{-1/2}),\quad n\in\mathbb{N}.$
It is easy to see that for $\lambda\in\mathbb{C}$ and $n\in\mathbb{N}$, $p(n)^{-1/2}p(n+1)p(n)^{-1/2}-\lambda I$ is invertible if and only if so is $p(n+1)-\lambda p(n)$. In turn, $p(n+1)-\lambda p(n)$ is invertible if and only if so is $p(n+1)p(n)^{-1}-\lambda I$. Hence,
$r(p(n)^{-1/2}p(n+1)p(n)^{-1/2})=r(p(n+1)p(n)^{-1}),\quad n\in\mathbb{N},$
and the proof is complete. ∎
The following example shows that in general the condition (iii) of Theorem 3.4 cannot be removed even if $H$ is finite dimensional.
Example 3.6.
Define $p\in\mathbf{B}(\mathbb{C}^{3})_{3}[z]$ by the formula:
$p(z)=\begin{bmatrix}z^{3}+16z+1&\frac{1}{2}z^{2}&2z\\ \frac{1}{2}z^{2}&\frac{1}{4}z+1&0\\ 2z&0&1\end{bmatrix},\quad z\in\mathbb{C}.$
An easy computation shows that for every $n\in\mathbb{N}$, $p(n)$ is self-adjoint and all leading principal minors of $p(n)$ are positive. By the Sylvester criterion (see [12, Theorem 7.2.5]), $p(n)$ is positive for every $n\in\mathbb{N}$. It follows that $p(n)$ is invertible. Obviously, $p(0)=I$. Now observe that (cf. [20, Eq. (1)] and the proof of Observation 3.5)
(3.8) $r(p(n+1)p(n)^{-1})$ $=r(p(n)^{-1/2}p(n+1)p(n)^{-1/2})$
$\leqslant\mathrm{tr}(p(n)^{-1/2}p(n+1)p(n)^{-1/2})$
$\leqslant 3r(p(n)^{-1/2}p(n+1)p(n)^{-1/2}),$
$=3r(p(n+1)p(n)^{-1}),\quad n\in\mathbb{N}.$
By the properties of a trace of matrix,
$\mathrm{tr}(p(n)^{-1/2}p(n+1)p(n)^{-1/2})=\mathrm{tr}(p(n+1)p(n)^{-1}),\quad n\in\mathbb{N}.$
Simple calculations gives us that
$\mathrm{tr}(p(n+1)p(n)^{-1})=\frac{n^{2}+208n+81}{4+65n},\quad n\in\mathbb{N}.$
Combining the above with (3.8), we get that
$r(p(n+1)p(n)^{-1})\xrightarrow{n\to\infty}\infty.$
By Observation 3.5,
$\sup\left\{\frac{\langle p(n+1)h,h\rangle}{\langle p(n)h,h\rangle}\colon n\in\mathbb{N},h\in H,h\not=0\right\}=\infty.$
However, under further assumption that the coefficients of $p$ are commuting, the condition (iii) of Theorem 3.4 is automatically satisfied.
Lemma 3.7.
Let $H=\mathbb{C}^{N}$ ($N\in\mathbb{Z}_{1}$) and let $p\in\mathbf{B}(H)_{m-1}[z]$ be a polynomial satisfying Theorem 3.4(i)–(ii). If the coefficients of $p$ are commuting, then the condition (iii) of Theorem 3.4 holds.
Proof.
Write $p(z)=A_{m-1}z^{m-1}+\ldots+A_{1}z+I$. From Observation 3.2 it follows that $A_{m-1},\ldots,A_{1}$ are self-adjoint. By [12, Theorem 2.5.5] there is no loss in generality in assuming that $A_{m-1},\ldots,A_{1}$ are diagonal matrices. This implies that $p$ is of the form
$p(z)=\begin{bmatrix}p_{1}(z)&0&\ldots&0\\ 0&p_{2}(z)&\ldots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\ldots&p_{N}(z)\end{bmatrix},\quad z\in\mathbb{C},$
where $p_{j}\in\mathbb{R}_{m-1}[z]$ are such that $p_{j}(n)>0$ ($n\in\mathbb{N},j=1,\ldots,N$). We have
$p(n)^{-1/2}p(n+1)p(n)^{-1/2}=\begin{bmatrix}\frac{p_{1}(n+1)}{p_{1}(n)}&0&\ldots&0\\ 0&\frac{p_{2}(n+1)}{p_{2}(n)}&\ldots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\ldots&\frac{p_{N}(n+1)}{p_{N}(n)}\end{bmatrix},\quad n\in\mathbb{N}.$
Hence,
$r(p(n)^{-1/2}p(n+1)p(n)^{-1/2})$ $\leqslant\mathrm{tr}(p(n)^{-1/2}p(n+1)p(n)^{-1/2})$
$=\sum_{j=1}^{N}\frac{p_{j}(n+1)}{p_{j}(n)}\xrightarrow{n\to\infty}N.$
By Observation 3.5, the condition Theorem 3.4(iii) is satisfied. ∎
If $H$ is infinite dimensional, then the conclusion of Lemma 3.7 does not hold.
Example 3.8.
For $k\in\mathbb{Z}_{1}$ define a polynomial $q_{k}\in\mathbb{R}_{2}[z]$ by the formula
$q_{k}(z)=\left(k+\frac{1}{k+1}\right)^{-1}\left(k+\frac{2}{3}\right)^{-1}\left(z-k-\frac{1}{k+1}\right)\left(z-k-\frac{2}{3}\right),\quad z\in\mathbb{C}.$
Let $H$ be an infinite dimensional separable Hilbert space and denote by $(e_{j})_{j\in\mathbb{Z}_{1}}\subset H$ its orthonormal basis. For $z\in\mathbb{C}$ define a diagonal operator $p(z)\in\mathbf{B}(H)$ as follows:
$p(z)e_{k}=q_{k}(z)e_{k},\quad k\in\mathbb{Z}_{1}.$
Clearly, $p\in\mathbf{B}(H)_{2}[z]$. Since $q_{k}(n)>0$ for every $n\in\mathbb{N}$, $p(n)$ is a positive operator. It is easy to see that $p(0)=I$. For every $n\in\mathbb{N}$ we have
$\langle p(n)e_{k},e_{k}\rangle=q_{k}(n)=\left(\frac{k+1}{k^{2}+k+1}n-1\right)\left(\frac{3}{3k+2}n-1\right)\xrightarrow{k\to\infty}1.$
From the above equality we deduce that for every $n\in\mathbb{N}$ there exists a constant $c_{n}\in(0,\infty)$ such that $\lVert p(n)h\rVert\geqslant c_{n}\lVert h\rVert$, $h\in H$. This implies that $\mathcal{R}(p(n))$ is closed for every $n\in\mathbb{N}$. Since $p(n)$ is positive, it follows from (2.1) that $p(n)$ is invertible for each $n\in\mathbb{N}$. However,
$\frac{\langle p(n+1)e_{n},e_{n}\rangle}{\langle p(n)e_{n},e_{n}\rangle}=\frac{\frac{1}{3}\left(1-\frac{1}{n+1}\right)}{\frac{1}{n+1}\cdot\frac{2}{3}}=\frac{n}{2}\xrightarrow{n\to\infty}\infty.$
Therefore, the condition (iii) in Theorem 3.4 is not satisfied.
As an another example of use of Theorem 3.1 we prove that every sequence $A_{1},\ldots,A_{m}\in\mathbf{B}(H)$ of invertible operators can be completed to the sequence of weights of $(m+2)$-isometric unilateral weighted shift. Abdullah and Le proved such a result for classical unilateral weighted shifts with scalar weights (see [1, Proposition 7]); the idea of the proof for shifts with operator weights is similar. Before we state the result, we introduce some notation. For distinct complex numbers $x_{0},\ldots,x_{m}\in\mathbb{C}$ we define the polynomials $\ell_{j}\in\mathbb{C}[x]$ ($j=0,\ldots,m$) as follows [^2] We stick to the convention that $\prod\varnothing=1$.:
$\ell_{j}(z)=\prod_{\begin{subarray}{c}k=0,\ldots,m\\ k\not=j\end{subarray}}\frac{z-x_{k}}{x_{j}-x_{k}},\quad z\in\mathbb{C}.$
It is a matter of routine to verify that $\ell_{j}(x_{j})=1$ for $j=0,\ldots,m$ and $\ell_{j}(x_{k})=0$ for $j,k\in\{0,\ldots,m\}$, $j\not=k$.
Theorem 3.9.
Let $H$ be a Hilbert space and let $A_{1},\ldots,A_{m}\in\mathbf{B}(H)$ be invertible operators. Then there exists an $(m+2)$-isometric unilateral weighted shift $S\in\mathbf{B}(\ell^{2}(\mathbb{N},H))$ with weights $(S_{j})_{j\in\mathbb{Z}_{1}}$ such that $S_{j}=A_{j}$ for every $j=1,\ldots,m$ and that $S_{j}$ is positive for $j\in\mathbb{Z}_{m+1}$.
Proof.
Let $\ell_{j}$ ($j=0,\ldots,m$) be the polynomials defined as above for the numbers $x_{j}=j$. For $h\in H$ set
$p(z):=\sum_{j=0}^{m}\widetilde{A}_{j}\ell_{j}(z),\quad z\in\mathbb{C},$
where $\widetilde{A}_{0}=I$ and $\widetilde{A}_{j}=\lvert A_{j}\cdots A_{1}\rvert^{2}$ for $j=1,\ldots,m$. Then $p\in\mathbf{B}(H)_{m}[z]$. Define $q\in\mathbf{B}(H)_{m+1}[z]$ by the formula
$q(z)=\alpha z(z-1)\cdots(z-m)I+p(z),\quad z\in\mathbb{C},h\in H,$
where [^3] Since $\sum_{j=0}^{m}\ell_{j}(n)$ is the polynomial in $n$ of degree $m$ and $n(n-1)\cdots(n-m)$ is the polynomial in $n$ of degree $m+1$, the supremum on the right hand side is finite.
(3.9) $\alpha>\sup_{n\geqslant m+1}\frac{\max_{j=0,\ldots,m}\lVert\widetilde{A}_{j}\rVert\sum_{j=0}^{m}\lvert\ell_{j}(n)\rvert}{n(n-1)\cdots(n-m)}.$
We have $q(0)=p(0)=I$ and $q(n)=p(n)=\lvert A_{n}\cdots A_{1}\rvert^{2}$ for $n=1,\ldots,m$. We will prove that $q(n)$ is positive for $n\in\mathbb{Z}_{m+1}$. If $n\in\mathbb{Z}_{m+1}$ and $h\in H$ is of norm one, then
$\langle q(n)h,h\rangle$ $=\alpha n(n-1)\cdots(n-m)+\sum_{j=0}^{m}\langle\widetilde{A}_{j}h,h\rangle\ell_{j}(n)$
$\geqslant\alpha n(n-1)\cdots(n-m)-\left\lvert\sum_{j=0}^{m}\langle\widetilde{A}_{j}h,h\rangle\ell_{j}(n)\right\rvert$
$\geqslant\alpha n(n-1)\cdots(n-m)-\sum_{j=0}^{m}\lvert\langle\widetilde{A_{j}}h,h\rangle\rvert\lvert\ell_{j}(n)\rvert$
$\geqslant\alpha n(n-1)\cdots(n-m)-\sum_{j=0}^{m}\lVert\widetilde{A_{j}}\rVert\lvert\ell_{j}(n)\rvert$
$\geqslant\alpha n(n-1)\cdots(n-m)-\max_{j=0,\ldots,m}\lVert\widetilde{A_{j}}\rVert\sum_{j=0}^{m}\lvert\ell_{j}(n)\rvert.$
By (3.9),
(3.10) $\langle q(n)h,h\rangle\geqslant\alpha n(n-1)\cdots(n-m)-\max_{j=0,\ldots,m}\lVert\widetilde{A_{j}}\rVert\sum_{j=0}^{m}\lvert\ell_{j}(n)\rvert>0$
From the above we also deduce that for every $n\in\mathbb{Z}_{m+1}$ there exists $c_{n}\in(0,\infty)$ such that $\langle q(n)h,h\rangle\geqslant c_{n}\lVert h\rVert^{2}$ for every $h\in H$. Therefore, $\mathcal{R}(q(n))$ is closed and, consequently, by (2.1), $q(n)$ is invertible for every $n\in\mathbb{Z}_{m+1}$. We define the sequence $(S_{j})_{j\in\mathbb{Z}_{1}}$ of invertible operators by induction. For $j=1,\ldots,m$ set $S_{j}=A_{j}$. If $S_{1},\ldots,S_{n}\in\mathbf{B}(H)$ ($n\in\mathbb{Z}_{m}$) are defined, then we set
$S_{n+1}:=\left[(S_{n}^{\ast})^{-1}\cdots(S_{1}^{\ast})^{-1}q(n+1)S_{n}^{-1}\cdots S_{1}^{-1}\right]^{1/2}.$
Clearly, each $S_{n}$ ($n\geqslant m+1$) is positive and invertible. Moreover,
(3.11) $\langle\lvert S_{n}\cdots S_{1}\rvert^{2}h,h\rangle=\langle q(n)h,h\rangle,\quad n\in\mathbb{Z}_{1},h\in H.$
It remains to check that $(S_{j})_{j\in\mathbb{Z}_{1}}$ is uniformly bounded. Note that for every $n\in\mathbb{Z}_{m+1}$ and $h\in H$ of norm one, we have
$\frac{\lVert S_{n+1}\cdots S_{n}h\rVert^{2}}{\lVert S_{n}\cdots S_{n}h\rVert^{2}}=\frac{\langle q(n+1)h,h\rangle}{\langle q(n)h,h\rangle}$
$=\frac{\alpha(n+1)n\cdots(n+1-m)+\sum_{j=0}^{m}\langle\widetilde{A_{j}}h,h\rangle\ell_{j}(n+1)}{\alpha n(n-1)\cdots(n-m)+\sum_{j=0}^{m}\langle\widetilde{A_{j}}h,h\rangle\ell_{j}(n)}$
(3.10)${{\leqslant}}\frac{\alpha(n+1)n\cdots(n+1-m)+\max_{j=0,\ldots,m}\lVert\widetilde{A_{j}}\rVert\sum_{j=0}^{m}\ell_{j}(n+1)}{\alpha n(n-1)\cdots(n-m)-\max_{j=0,\ldots,m}\lVert\widetilde{A_{j}}\rVert\sum_{j=0}^{m}\lvert\ell_{j}(n)\rvert}.$
Since the polynomial $\ell_{j}$ is of degree $m$ for every $j=0,1\ldots,m$, it is a matter of routine to verify that
$\lim_{n\to\infty}\frac{\alpha(n+1)n\cdots(n+1-m)+\max_{j=0,\ldots,m}\lVert\widetilde{A_{j}}\rVert\sum_{j=0}^{m}\ell_{j}(n+1)}{\alpha n(n-1)\cdots(n-m)-\max_{j=0,\ldots,m}\lVert\widetilde{A_{j}}\rVert\sum_{j=0}^{m}\lvert\ell_{j}(n)\rvert}=1.$
It follows that there exists $C\in(0,\infty)$ such that for every $n\in\mathbb{Z}_{m+1}$,
$\frac{\alpha(n+1)n\cdots(n+1-m)+\max_{j=0,\ldots,m}\lVert\widetilde{A_{j}}\rVert\sum_{j=0}^{m}\ell_{j}(n+1)}{\alpha n(n-1)\cdots(n-m)-\max_{j=0,\ldots,m}\lVert\widetilde{A_{j}}\rVert\sum_{j=0}^{m}\lvert\ell_{j}(n)\rvert}\leqslant C.$
Thus,
$\lVert S_{n+1}\cdots S_{n}h\rVert^{2}\leqslant C\lVert S_{n}\cdots S_{n}h\rVert^{2},\quad n\in\mathbb{Z}_{m+1},h\in H.$
Since $\mathcal{R}(S_{n}\cdots S_{1})=H$, we have
$\lVert S_{n+1}\rVert\leqslant\sqrt{C},\quad n\in\mathbb{Z}_{m+1}.$
Therefore,
$\sup_{n\in\mathbb{Z}_{1}}\lVert S_{n}\rVert\leqslant\max\{\sqrt{C},\lVert S_{1}\rVert,\ldots,\lVert S_{m+1}\rVert\}.$
Using (3.11) we deduce from Theorem 3.1 that the weighted shift $S\in\mathbf{B}(\ell^{2}(\mathbb{Z},H))$ with weights $(S_{j})_{j\in\mathbb{Z}_{1}}$ is $(m+2)$-isometric and such that $S_{j}=A_{j}$ for every $j=1,\ldots,m$. ∎
Theorem 3.9 states (in particular) that if $A_{1}\in\mathbf{B}(H)$ is invertible, then there exists a 3-isometric unilateral weighted shift $S\in\mathbf{B}(\ell^{2}(\mathbb{N},H))$ with weights starting from $A_{1}$. However, it may not be possible to find such a shift, which is 2-isometric. It was proved by Jabłoński in [13, Proposition 3.2 and Remark 3.4] that $A_{1}$ is a first weight of 2-isometric unilateral weighted shifts if and only if $A_{1}^{\ast}A_{1}\geqslant I$.
Next, let us make some remarks about commutativity of weights of $m$-isometric unilateral weighted shift. We start from general result.
Theorem 3.10.
Let $H$ be a Hilbert space and let $S\in\mathbf{B}(\ell^{2}(\mathbb{N},H))$ be an $m$-isometric unilateral weighted shifts with positive and invertible weights $(S_{j})_{j\in\mathbb{Z}_{1}}\subset\mathbf{B}(H)$. The following conditions are equivalent:
  1. (i)
    all weights of $S$ commute,
  2. (ii)
    $S_{1},\ldots,S_{m-1}$ commute,
  3. (iii)
    if $p(z)=A_{m-1}z^{m-1}+\ldots+A_{1}z+I\in\mathbf{B}(H)_{m-1}[z]$ is a polynomial as in Theorem 3.1(iii), then $A_{m-1},\ldots,A_{1}$ commute.
In the proof we use the following lemma.
Lemma 3.11.
Let $H$ be a Hilbert space and let $(S_{j})_{n\in\mathbb{Z}_{1}}\subset\mathbf{B}(H)$ be a sequence of positive and invertible operators. Assume that for every $n,m\in\mathbb{Z}_{1}$, $\lvert S_{n}\cdots S_{1}\rvert$ commutes with $\lvert S_{m}\cdots S_{1}\rvert$. Then $S_{n}$ commutes with $S_{m}$ for every $n,m\in\mathbb{Z}_{1}$.
Proof.
We prove by induction that for every $n\in\mathbb{Z}_{1}$ the operators $S_{1},\ldots,S_{n}$ are commuting. First consider the case $n=2$. By the assumption, we have
$S_{1}S_{1}S_{2}S_{2}S_{1}=S_{1}\lvert S_{2}S_{1}\rvert^{2}=\lvert S_{2}S_{1}\rvert^{2}S_{1}=S_{1}S_{2}S_{2}S_{1}S_{1}.$
Since $S_{1}$ is invertible, it follows that
$S_{1}S_{2}^{2}=S_{2}^{2}S_{1}.$
By the square root theorem (see [22, p.265]), $S_{1}$ commutes with $S_{2}$. Now suppose that for some $n\in\mathbb{Z}_{2}$, $S_{1},\ldots,S_{n}$ commute. We will show that
(3.12) $S_{k}\lvert S_{n+1}\cdots S_{1}\rvert=\lvert S_{n+1}\cdots S_{1}\rvert S_{k},\quad k=1,\ldots,n.$
For $k=1$, the above equality follows directly from the assumption. Let $2\leqslant k\leqslant n$. It is enough to prove that
(3.13) $S_{k}^{-2}\lvert S_{n+1}\cdots S_{1}\rvert^{2}=\lvert S_{n+1}\cdots S_{1}\rvert^{2}S_{k}^{-2}.$
Indeed, (3.13) implies that
$S_{k}^{2}\lvert S_{n+1}\cdots S_{1}\rvert^{2}=\lvert S_{n+1}\cdots S_{1}\rvert^{2}S_{k}^{2}.$
Using the square root theorem we obtain (3.12). For the proof of (3.13) observe that, by [5, Lemma 2.1] and by the inductive hypothesis, we have
$S_{k}^{-2}\lvert S_{n+1}\cdots S_{1}\rvert^{2}S_{k-1}^{-2}\cdots S_{1}^{-2}$ $=S_{k}^{-2}\lvert S_{n+1}\cdots S_{1}\rvert^{2}\lvert S_{k-1}\cdots S_{1}\rvert^{-2}$
$=S_{k}^{-2}\lvert S_{k-1}\cdots S_{1}\rvert^{-2}\lvert S_{n+1}\cdots S_{1}\rvert^{2}$
$=S_{k}^{-2}S_{k-1}^{-2}\cdots S_{1}^{-2}\lvert S_{n+1}\cdots S_{1}\rvert^{2}$
$=\lvert S_{k}\cdots S_{1}\rvert^{-2}\lvert S_{n+1}\cdots S_{1}\rvert^{2}$
$=\lvert S_{n+1}\cdots S_{1}\rvert^{2}\lvert S_{k}\cdots S_{1}\rvert^{-2}$
$=\lvert S_{n+1}\cdots S_{1}\rvert^{2}S_{k}^{-2}S_{k-1}^{-2}\cdots S_{1}^{-2}.$
Since $\mathcal{R}(S_{k-1}^{-2}\cdots S_{1}^{-1})=H$, (3.13) follows from the above. Now, again by [5, Lemma 2.1],
$\lvert S_{n}\cdots S_{1}\rvert^{-1}\lvert S_{n+1}\cdots S_{1}\rvert^{2}\lvert S_{n}\cdots S_{1}\rvert^{-1}=S_{1}^{-1}\ldots S_{n}^{-1}\lvert S_{n+1}\cdots S_{1}\rvert^{2}S_{1}^{-1}\ldots S_{n}^{-1}$
$=S_{n}^{-1}\ldots S_{1}^{-1}S_{1}\ldots S_{n}S_{n+1}^{2}S_{n}\ldots S_{1}S_{1}^{-1}\ldots S_{n}^{-1}=S_{n+1}^{2}.$
By (3.12), we have
$S_{k}S_{n+1}^{2}$ $=S_{k}\lvert S_{n}\cdots S_{1}\rvert^{-1}\lvert S_{n+1}\cdots S_{1}\rvert^{2}\lvert S_{n}\cdots S_{1}\rvert^{-1}$
$=\lvert S_{n}\cdots S_{1}\rvert^{-1}\lvert S_{n+1}\cdots S_{1}\rvert^{2}\lvert S_{n}\cdots S_{1}\rvert^{-1}S_{k}$
$=S_{n+1}^{2}S_{k}.$
By the square root theorem, $S_{k}S_{n+1}=S_{n+1}S_{k}$. This completes the proof. ∎
Proof of Theorem 3.10.
The implication (i)$\Longrightarrow$(ii) obviously holds; the implication (ii)$\Longrightarrow$(iii) follows directly from Lemma 3.2. For the proof of implication (iii)$\Longrightarrow$(i) note that for every $n,m\in\mathbb{Z}_{1}$, by (iii), we have
$\lvert S_{n}\cdots S_{1}\rvert^{2}\lvert S_{n}\cdots S_{1}\rvert^{2}=p(n)p(m)=p(m)(n)=\lvert S_{m}\cdots S_{1}\rvert^{2}\lvert S_{n}\cdots S_{1}\rvert^{2}.$
The application of Lemma 3.11 completes the proof. ∎
Corollary 3.12.
Let $H$ be a Hilbert space and let $S\in\mathbf{B}(\ell^{2}(\mathbb{N},H))$ be an $m$-isometric unilateral weighted shifts with positive and invertible weights such that the polynomial $p\in\mathbf{B}_{m-1}(H)[z]$ as in Theorem 3.1(iii) is of the form $p(z)=Az^{m-1}+I$ for some $A\in\mathbf{B}(H)$. Then the weights of $S$ commute. In particular, if $S$ is 2-isometric, then the weights of $S$ commute.
The fact that the weights of 2-isometric unilateral weighted shifts with positive and invertible weights are commuting was pointed out by Jabłoński in [13, Theorem 3.3]; however, the proof is different from ours.
[…]
The following example show that for every $m\in\mathbb{Z}_{3}$ we can find $m$-isometric weighted shift with positive, invertible and non-commuting weights.
Example 3.13.
Let $H$ be a Hilbert space of dimension greater than one. Assume that $A_{1},\ldots,A_{m}\subset\mathbf{B}(H)$ ($m\in\mathbb{Z}_{2}$) are positive and invertible operators and at least two of them do not commute. Set $A_{0}=I$. Define a polynomial $p\in\mathbf{B}(H)_{m}[z]$ by the formula:
$p(z)=\sum_{j=0}^{m}A_{j}z^{j},\quad z\in\mathbb{C}.$
It is easy to see that $p(n)$ is positive and invertible for every $n\in\mathbb{Z}_{1}$ and that $p(0)=I$. Since each $A_{j}$ is positive and invertible, there exist $c,C\in(0,\infty)$ such that
$C\lVert h\rVert^{2}\geqslant\langle A_{j}h,h\rangle=\lVert A_{j}^{1/2}h\rVert^{2}\geqslant c\lVert h\rVert^{2},\quad h\in H,j=0,1,\ldots,m.$
Hence,
$\frac{\langle p(n+1)h,h\rangle}{\langle p(n)h,h\rangle}\leqslant\frac{C\sum_{j=0}^{m}(n+1)^{j}}{c\sum_{j=0}^{m}n^{j}},\quad n\in\mathbb{N},h\in H,h\not=0.$
Since
$\frac{C\sum_{j=0}^{m}(n+1)^{j}}{c\sum_{j=0}^{m}n^{j}}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}\frac{C}{c},$
the condition (iii) of Theorem 3.4 is satisfied. By Theorem 3.4 there exists $(m+1)$-isometric weighted shift $S\in\mathbf{B}(\ell^{2}(\mathbb{Z},H))$ with positive and invertible weights $(S_{j})_{j\in\mathbb{Z}_{1}}$ such that
$\lvert S_{n}\cdots S_{1}\rvert^{2}=\sum_{j=0}^{m}A_{j}n^{j},\qquad n\in\mathbb{N}.$
By Theorem 3.10, the weights of $S$ do not commute.
4. $m$-isometric bilateral weighted shifts
The aim of this section is to provide counterpart of results presented in Section 3 for bilateral weighted shifts. Let us introduce some notation. If $H$ is a Hilbert space, then for $k\in\mathbb{Z}$ we denote
$L_{k}=\{(h_{j})_{j\in\mathbb{Z}}\in\ell^{2}(\mathbb{Z},H)\colon h_{j}=0\text{ for }j<k\}.$
If $S\in\mathbf{B}(\ell^{2}(\mathbb{Z},H))$ is a bilateral weighted shift with weights $(S_{j})_{j\in\mathbb{Z}}\subset\mathbf{B}(H)$, then by $S_{k}$ we denote the restriction of $S$ to the subspace $L_{k}$. Below we gather basic properties of the operators $S_{k}$; we leave the proof to the reader.
Observation 4.1.
Let $H$ be a Hilbert space and let $S\in\mathbf{B}(\ell^{2}(\mathbb{Z},H))$ be a bilateral weighted shift with weights $(S_{j})_{j\in\mathbb{Z}}\subset\mathbf{B}(H)$. Then:
  1. (i)
    for every $k\in\mathbb{Z}$, the subspace $L_{k}$ is an invariant subspace for $S$ and for $S^{\ast n}S^{n}$ ($n\in\mathbb{Z}_{1}$),
  2. (ii)
    for every $k\in\mathbb{Z}$, $S_{k}$ is unitarily equivalent to the unilateral weighted shift with weights $(S_{j+k})_{j\in\mathbb{Z}_{1}}$,
  3. (iii)
    for $m\in\mathbb{Z}_{1}$, $S$ is $m$-isometric if and only if $S_{k}$ is $m$-isometric for every $k\in\mathbb{Z}$.
Now we are in the position to state and prove the main result of this section.
Theorem 4.2.
Let $H$ be a Hilbert space and let $S\in\mathbf{B}(\ell^{2}(\mathbb{Z},H))$ be a bilateral weighted shift with invertible weights $(S_{i})_{i\in\mathbb{Z}}\subset\mathbf{B}(H)$. For $m\in\mathbb{Z}_{1}$ the following conditions are equivalent:
  1. (i)
    $S$ is $m$-isometric,
  2. (ii)
    there exists $p\in\mathbf{B}(H)_{m-1}[z]$ such that $p(0)=I$ and
    (4.1) $p(n)$ $=\lvert S_{n}\cdots S_{1}\rvert^{2},\quad n\in\mathbb{Z}_{1},$
    (4.2) $p(-n)$ $=\lvert S_{-n+1}^{\ast}\cdots S_{0}^{\ast}\rvert^{-2},\quad n\in\mathbb{Z}_{1}.$
Moreover, if (ii) holds, then
(4.3) $\lVert S\rVert^{2}=\sup\left\{\frac{\langle p(n+1)h,h\rangle}{\langle p(n)h,h\rangle}\colon n\in\mathbb{Z},h\in H,h\not=0\right\}.$
Proof.
(i)$\Longrightarrow$(ii). By Observation 4.1 and Theorem 3.1, for every $k\in\mathbb{Z}$, $k\leqslant 0$, there exists $p_{k}\in\mathbf{B}(H)_{m-1}[z]$ such that $p_{k}(0)=I$ and
(4.4) $p_{k}(n)=\lvert S_{k+n}\cdots S_{k+1}\rvert^{2},\quad n\in\mathbb{Z}_{1}.$
Observe that
$p_{k}(n-k)=\lvert S_{n}\cdots S_{k+1}\rvert^{2}$ $=S_{k+1}^{\ast}\cdots S_{0}^{\ast}\lvert S_{n}\cdots S_{1}\rvert^{2}S_{0}\cdots S_{k+1},$
$=S_{k+1}^{\ast}\cdots S_{0}^{\ast}p_{0}(n)S_{0}\cdots S_{k+1}\quad n\in\mathbb{Z}_{1},k\in\mathbb{Z},k<0.$
Hence, we have
$p_{0}(n)=(S_{k+1}^{\ast}\cdots S_{0}^{\ast})^{-1}p_{k}(n-k)(S_{0}\cdots S_{k+1})^{-1},\quad n\in\mathbb{Z}_{1},k\in\mathbb{Z},k<0.$
Since
$(S_{k+1}^{\ast}\cdots S_{0}^{\ast})^{-1}p_{k}(z-k)(S_{0}\cdots S_{k+1})^{-1}\in\mathbf{B}(H)_{m-1}[z],\quad k\in\mathbb{Z},k<0,$
it follows that
$(S_{k+1}^{\ast}\cdots S_{0}^{\ast})^{-1}p_{k}(z-k)(S_{0}\cdots S_{k+1})^{-1}=p_{0}(z),\quad z\in\mathbb{C},k\in\mathbb{Z},k<0.$
In particular,
$p_{0}(k)$ $=(S_{k+1}^{\ast}\cdots S_{0}^{\ast})^{-1}p_{k}(0)(S_{0}\cdots S_{k+1})^{-1}$
$=(S_{k+1}^{\ast}\cdots S_{0}^{\ast})^{-1}(S_{0}\cdots S_{k+1})^{-1}$
$=(S_{0}\cdots S_{k+1}S_{k+1}^{\ast}\cdots S_{0}^{\ast})^{-1}$
$=\lvert S_{k+1}^{\ast}\cdots S_{0}\rvert^{-2},\quad k\in\mathbb{Z},k<0.$
Therefore, $p_{0}$ is the polynomial satisfying (ii).
(ii)$\Longrightarrow$(i). Arguing as in the proof of the converse implication, we can easily verify that for every $k\in\mathbb{Z}$, $k<0$, the polynomial $p_{k}\in\mathbf{B}(H)_{m-1}[z]$ given by
(4.5) $p_{k}(z)=S_{k+1}^{\ast}\cdots S_{0}^{\ast}p(z+k)S_{0}\cdots S_{k+1},\quad z\in\mathbb{C},$
satisfies Theorem 3.1(iii) for $S_{k}$. Moreover, $p$ satisfies Theorem 3.1 for $S_{0}$. Hence, $S_{k}$ is $m$-isometric for every $k\in\mathbb{Z}$, $k\leqslant 0$. Since $S_{k}\subset S_{0}$ for every $k\in\mathbb{Z}_{1}$, $S_{k}$ is $m$-isometric also for $k\in\mathbb{Z}_{1}$. The application of Observation 4.1(iii) gives (i).
Now we prove the ”moreover” part. Proceeding as in the proof of Theorem 3.1 we obtain that
$\lVert S_{n+1}\rVert^{2}=\sup\left\{\frac{\langle p(n+1)h,h\rangle}{\langle p(n)h,h\rangle}\colon h\in H,h\not=0\right\},\quad n\in\mathbb{N}.$
Next, note that
$\lvert S_{-n+1}^{\ast}\cdots S_{0}^{\ast}\rvert^{-2}=S_{0}^{\ast-1}\cdots S_{-n+1}^{\ast-1}S_{-n+1}^{-1}\cdots S_{0}^{-1},\quad n\in\mathbb{Z}_{1}$
Hence, if $n\in\mathbb{Z}_{2}$, then, by the fact that $\mathcal{R}(S_{-n+1}^{-1}\cdots S_{0}^{-1})=H$, we have
$\lVert S_{-n+1}\rVert^{2}$ =$\sup\left\{\frac{\lVert S_{-n+1}h\rVert^{2}}{\lVert h\rVert^{2}}\colon h\in H,h\not=0\right\}$
=$\sup\left\{\frac{\lVert S_{-n+1}S_{-n+1}^{-1}\cdots S_{0}^{-1}h\rVert^{2}}{\lVert S_{-n+1}^{-1}\cdots S_{0}^{-1}h\rVert^{2}}\colon h\in H,h\not=0\right\}$
=$\sup\left\{\frac{\langle S_{0}^{\ast-1}\cdots S_{-n+2}^{\ast-1}S_{-n+2}^{-1}\cdots S_{0}^{-1}h,h\rangle}{\langle S_{0}^{\ast-1}\cdots S_{-n+1}^{\ast-1}S_{-n+1}^{-1}\cdots S_{0}^{-1}h,h\rangle}\colon h\in H,h\not=0\right\}$
=$\sup\left\{\frac{\langle\lvert S_{-n+2}^{\ast}\cdots S_{0}^{\ast}\rvert^{-2}h,h\rangle}{\langle\lvert S_{-n+1}^{\ast}\cdots S_{0}^{\ast}\rvert^{-2}h,h\rangle}\colon h\in H,h\not=0\right\}$
=$\sup\left\{\frac{\langle p(-n+1)h,h\rangle}{\langle p(-n)h,h\rangle}\colon h\in H,h\not=0\right\}.$
By the fact that $p(0)=I$, arguing similarly as above we get that
$\lVert S_{0}\rVert^{2}=\left\{\frac{\langle p(0)h,h\rangle}{\langle p(-1)h,h\rangle}\colon h\in H,h\not=0\right\}.$
Therefore, (4.3) follows from (2.7). ∎
Below we state without the proof the counterpart of Corollary 3.3 for bilateral weighted shifts.
Corollary 4.3.
Let $H$ and $S$ be as in Theorem 4.2.
  1. (i)
    If $S$ is $m$-isometric and the polynomial $p\in\mathbf{B}(H)_{m-1}[z]$ is as in Theorem 4.2(ii), then $\lvert S_{n}\rvert^{2}$ is unitarily equivalent to $p(n-1)^{-1/2}p(n)p(n-1)^{-1/2}$ and $\lvert S_{-n+1}^{\ast}\rvert^{2}$ is unitarily equivalent to $p(-n+1)^{-1/2}p(-n)p(-n+1)^{-1/2}$ for every $n\in\mathbb{Z}_{1}$.
  2. (ii)
    If the weights of $S$ are doubly commuting, then $m$-isometricity of $S$ is equivalent to the existence of a polynomial $p\in\mathbf{B}(H)_{m-1}[z]$ satisfying $p(0)=I$ and
    $\lvert S_{n}\rvert^{2}$ $=p(n)p(n-1)^{-1}=p(n-1)^{-1}p(n),\quad n\in\mathbb{Z}_{1},$
    $\lvert S_{-n+1}^{\ast}\rvert^{2}$ $=p(-n)p(-n+1)^{-1}=p(-n+1)^{-1}p(-n),\quad n\in\mathbb{Z}_{1}.$
As in the case of unilateral weighted shifts, we have the procedure of constructing of all $m$-isometric bilateral weighted shifts with invertible and positive weights from a given polynomial $p\in\mathbf{B}(H)_{m-1}[z]$ satisfying certain conditions.
Theorem 4.4.
Let $H$ be a Hilbert space. If $p\in\mathbf{B}(H)_{m-1}[z]$ is a polynomial such that
  1. (i)
    $p(0)=I$,
  2. (ii)
    $p(n)$ is positive and invertible for every $n\in\mathbb{Z}$,
  3. (iii)
    $D(p):=\sup\left\{\frac{\langle p(n+1)h,h\rangle}{\langle p(n)h,h\rangle}\colon n\in\mathbb{Z},h\in H,h\not=0\right\}<\infty$,
then there exists an $m$-isometric bilateral weighted shift with positive and invertible weights $(S_{j})_{j\in\mathbb{Z}}$ satisfying (4.1) and (4.2). Moreover, $\lVert S\rVert=\sqrt{D(p)}$.
Proof.
Define the sequence $(S_{j})_{j\in\mathbb{Z}}\subset\mathbf{B}(H)$ satisfying (4.1) and (4.2) by induction. Set $S_{1}:=p(1)^{1/2}$ and $S_{0}:=p(-1)^{-1/2}$. Clearly, $S_{1}$ and $S_{0}$ are positive and invertible and (4.1)-(4.2) hold for $n=-1,0,1$. Now assume that for some $k\in\mathbb{Z}_{1}$, $S_{-k+1},\ldots,S_{0},S_{1},\ldots,S_{k}$ are defined to be positive and invertible operators satisfying (4.1)–(4.2) for $n=-k,\ldots,k$. Then
$\widetilde{S}_{k+1}$ $:=S_{1}^{-1}\cdots S_{k}^{-1}p(k+1)S_{k}^{-1}\cdots S_{1}^{-1},$
$\widetilde{S}_{-k}$ $:=S_{-k+1}^{-1}\cdots S_{0}^{-1}p(-k-1)S_{0}^{-1}\cdots S_{-k+1}^{-1}$
are positive and invertible and so are $S_{k+1}:=\widetilde{S}_{k+1}^{1/2}$ and $S_{-k}:=\widetilde{S}_{-k}^{-1/2}$. Proceeding as in the proof of Theorem 3.4 we check that (4.1) holds for $n=k+1$. To verify that (4.2) holds for $n=-k-1$ observe that
$\langle\lvert S_{-k}\cdots S_{0}\rvert^{-2}h,h\rangle$ $=\langle S_{0}^{-1}\cdots S_{-k+1}^{-1}\cdot\widetilde{S}_{-k}\cdot S_{-k+1}^{-1}\cdots S_{0}^{-1}h,h\rangle$
$=\langle p(-k-1)h,h\rangle.$
We will show that the sequence $(S_{j})_{j\in\mathbb{Z}}$ defined above is uniformly bounded. Arguing as in the proof of Theorem 3.4, we have $\lVert S_{n}\rVert^{2}\leqslant D(p)$ for every $n\in\mathbb{Z}_{1}$. In turn, by (iii),
$\langle\lvert S_{-n+2}\cdots S_{0}\rvert^{-2}h,h\rangle$ $=\langle p(-n+1)h,h\rangle$
$\leqslant D(p)\langle p(-n)h,h\rangle=\langle\lvert S_{-n+1}\cdots S_{0}\rvert^{-2}h,h\rangle,\quad n\in\mathbb{Z}_{2}.$
Substituting $S_{0}\cdots S_{-n+1}h$ in place of $h$ in the above inequality and using the fact that
$S_{-n+1}\cdots S_{0}\lvert S_{-n+2}\cdots S_{0}\rvert^{-2}S_{0}\cdots S_{-n+1}=S_{-n+2}^{2}$
and
$S_{-n+1}\cdots S_{0}\lvert S_{-n+1}\cdots S_{0}\rvert^{-2}S_{0}\cdots S_{-n+1}=I,$
we obtain that
$\lVert S_{-n+1}h\rVert^{2}=\langle S_{-n+1}^{2}h,h\rangle\leqslant D(p)\langle h,h\rangle=D(p)\lVert h\rVert^{2},\quad h\in H,n\in\mathbb{Z}_{2},$
so $\lVert S_{-n+1}\rVert^{2}\leqslant D(p)$ for every $n\in\mathbb{Z}_{2}$. Similar argument shows that $\lVert S_{0}\rVert^{2}\leqslant D(p)$. We deduce from Theorem 4.2 that the weighted shift $S\in\mathbf{B}(\ell^{2}(\mathbb{Z},H))$ with weights $(S_{j})_{j\in\mathbb{Z}}$ is $m$-isometric and $\lVert S\rVert=\sqrt{D(p)}$. ∎
Remark 4.5.
We can prove, in essentially the same way as in the proof of Observation 3.5, that if $p\in\mathbf{B}_{m-1}(H)$ satisfies Theorem 4.4(i)–(ii), then
$D(p)=\sup\left\{r(p(n+1)p(n)^{-1})\colon n\in\mathbb{Z}\right\}.$
As regards commutativity of weights of $m$-isometric bilateral weighted shift, we have the following result (cf. Theorem 3.10).
Theorem 4.6.
Let $H$ be a Hilbert space and let $S\in\mathbf{B}(\ell^{2}(\mathbb{Z},H))$ be an $m$-isometric bilateral weighted shift with positive and invertible weights $(S_{j})_{j\in\mathbb{Z}}\subset\mathbf{B}(H)$. The following conditions are equivalent:
  1. (i)
    all weights of $S$ commute,
  2. (ii)
    $S_{1},\ldots,S_{m-1}$ commute,
  3. (iii)
    if $p(z)=A_{m-1}z^{m-1}+\ldots+A_{1}z+I\in\mathbf{B}(H)_{m-1}[z]$ satisfies Theorem 4.6(ii), then $A_{1},\ldots,A_{m-1}$ commute.
Proof.
The implications (i)$\Longrightarrow$(ii) and (ii)$\Longrightarrow$(iii) may be proved in exactly the same way as in Theorem 3.10.
(iii)$\Longrightarrow$(i). By Observation 4.1, $S_{k}$ is $m$-isometric for every $k\in\mathbb{Z}$, $k\leqslant 0$. It is sufficient to show that for every $k<0$ the weights of $S_{k}$ commute. Proceeding as in the proof of Theorem 4.2, we get that $p$ is the polynomial satisfying Theorem 3.1(iii) for $S_{0}$ and the polynomials $p_{k}$ ($k<0$) satisfying Theorem 3.1(iii) for $S_{k}$ are given by (4.5). By Theorem 3.10, it is enough to prove that the coefficients of $p_{k}$ commute. From (4.5) it follows that
$p_{k}(z)=\sum_{j=1}^{m-1}B_{k}^{\ast}A_{j}B_{k}(z+k)^{j}+B_{k}^{\ast}B_{k},\quad k\in\mathbb{Z},k<0,$
where $B_{k}=S_{0}\cdots S_{k+1}$ ($k<0$). By Lemma 3.2, for every $j=1,\ldots,m-1$, $A_{j}$ is a linear combination of operators $I,\lvert S_{1}\rvert^{2},\cdots,\lvert S_{m-1}\cdots S_{1}\rvert^{2}$ with real coefficients. Hence, it suffice to show that for every $j,\ell\in\left\{1,\ldots m-1\right\}$,
(4.6) $B_{k}^{\ast}\lvert S_{j}\cdots S_{1}\rvert^{2}B_{k}B_{k}^{\ast}\lvert S_{\ell}\cdots S_{1}\rvert^{2}B_{k}=B_{k}^{\ast}\lvert S_{\ell}\cdots S_{1}\rvert^{2}B_{k}B_{k}^{\ast}\lvert S_{j}\cdots S_{1}\rvert^{2}B_{k}$
and
(4.7) $B_{k}^{\ast}\lvert S_{j}\cdots S_{1}\rvert^{2}B_{k}B_{k}^{\ast}B_{k}=B_{k}^{\ast}B_{k}B_{k}^{\ast}\lvert S_{j}\cdots S_{1}\rvert^{2}B_{k}.$
Since $B_{k}$ is invertible, (4.6) and (4.7) are equivalent to
$\lvert S_{j}\cdots S_{1}\rvert^{2}B_{k}B_{k}^{\ast}\lvert S_{\ell}\cdots S_{1}\rvert^{2}=\lvert S_{\ell}\cdots S_{1}\rvert^{2}B_{k}B_{k}^{\ast}\lvert S_{j}\cdots S_{1}\rvert^{2}$
and
$\lvert S_{j}\cdots S_{1}\rvert^{2}B_{k}B_{k}^{\ast}=B_{k}B_{k}^{\ast}\lvert S_{j}\cdots S_{1}\rvert^{2},$
respectively. But, by the equality $B_{k}B_{k}^{\ast}=\lvert S_{k+1}\cdots S_{0}\rvert^{2}$, we deduce from (iii) that
$\lvert S_{j}\cdots S_{1}\rvert^{2}B_{k}B_{k}^{\ast}\lvert S_{\ell}\cdots S_{1}\rvert^{2}$ $=p(j)p(k)p(\ell)$
$=p(\ell)p(k)p(j)=\lvert S_{\ell}\cdots S_{1}\rvert^{2}B_{k}B_{k}^{\ast}\lvert S_{j}\cdots S_{1}\rvert^{2}$
and
$\lvert S_{j}\cdots S_{1}\rvert^{2}B_{k}B_{k}^{\ast}=p(j)p(k)=p(k)p(j)=B_{k}B_{k}^{\ast}\lvert S_{j}\cdots S_{1}\rvert^{2}$
for every $j,\ell\in\left\{1,\ldots,m-1\right\}$. Thus, (4.6) and (4.7) hold, so the coefficients of $p_{k}$ commute. By Theorem 3.10, the weights of $S_{k}$ commute for every $k\in\mathbb{Z}$, $k<0$. ∎
In [7, Problem 4.9] Bermúdez et.al. posed a question whether the adjoint of $m$-isometric invertible operator is also $m$-isometric. In the following result we characterize bilateral weighted shifts with operator weights, which possess this property.
Theorem 4.7.
Let $H$ be a Hilbert space and let $S$ be an $m$-isometric bilateral weighted shift with invertible weights $(S_{j})_{j\in\mathbb{Z}}$. The following conditions are equivalent:
  1. (i)
    $S^{\ast}$ is $m$-isometric,
  2. (ii)
    the polynomial $p\in\mathbf{B}(H)_{m-1}[z]$ as in Theorem 4.2(ii) is invertible in $\mathbf{B}(H)_{m-1}[z]$, that is, there exists a polynomial $\widetilde{p}\in\mathbf{B}(H)_{m-1}[z]$ such that
    (4.8) $p(z)\cdot\widetilde{p}(z)=\widetilde{p}(z)\cdot p(z)=I,\quad z\in\mathbb{C}.$
Proof.
By Lemma 2.1, $S^{\ast}$ is unitarily equivalent to the bilateral weighted shift $\widehat{S}\in\mathbf{B}(\ell^{2}(\mathbb{Z},H))$ with weights $(S_{-j+1}^{\ast})_{j\in\mathbb{Z}}$; $S^{\ast}$ is $m$-isometric if and only if so is $\widehat{S}$. By Theorem 4.2, $\widehat{S}$ is $m$-isometric if and only if there exists a polynomial $\widehat{p}\in\mathbf{B}(H)_{m-1}[z]$ such that
(4.9) $\widehat{p}(0)$ $=I,$
(4.10) $\widehat{p}(n)$ $=\lvert S_{-n+1}^{\ast}\cdots S_{0}^{\ast}\rvert^{2},\quad n\in\mathbb{Z}_{1},$
(4.11) $\widehat{p}(-n)$ $=\lvert S_{n}\cdots S_{1}\rvert^{-2},\quad n\in\mathbb{Z}_{1}.$
(i)$\Longrightarrow$(ii). Let $\widehat{p}\in\mathbf{B}(H)_{m-1}[z]$ satisfy (4.9)-(4.11) and define $\widetilde{p}(z)=\widehat{p}(-z)$, $z\in\mathbb{C}$. Then $\widetilde{p}\in\mathbf{B}(H)_{m-1}[z]$ and
$p(0)\widetilde{p}(0)$ $=p(0)\widehat{p}(0)=I\cdot I=I,$
$p(n)\widetilde{p}(n)$ $=p(n)\widehat{p}(-n)=\lvert S_{n}\cdots S_{1}\rvert^{2}\lvert S_{n}\cdots S_{1}\rvert^{-2}=I,\quad n\in\mathbb{Z}_{1},$
$p(-n)\widetilde{p}(-n)$ $=p(-n)\widehat{p}(n)=\lvert S_{-n+1}^{\ast}\cdots S_{0}^{\ast}\rvert^{-2}\lvert S_{-n+1}^{\ast}\cdots S_{0}^{\ast}\rvert^{2}=I,\quad n\in\mathbb{Z}_{1}.$
Since $p\cdot\widetilde{p}\in\mathbf{B}_{2m-2}[z]$ and $p(n)\widetilde{p}(n)=I$ for every $n\in\mathbb{Z}$, it follows that $p\cdot\widetilde{p}=I$. Similarly, $\widetilde{p}\cdot p=I$.
(ii)$\Longrightarrow$(i). Proceeding as in the proof of the converse implication, we get that the polynomial $\widehat{p}(z)=\widetilde{p}(-z)$, $z\in\mathbb{C}$, satisfies (4.9)-(4.11). ∎
As a corollary, we obtain that the only $m$-isometric classical bilateral weighted shifts (with scalar weights), for which the adjoint is also $m$-isometric, are unitary. This answers the question in [7, Problem 4.9] in negative.
Corollary 4.8.
Let $S\in\mathbf{B}(\ell^{2}(\mathbb{Z},\mathbb{C}))$ be an $m$-isometric bilateral weighted shift with weights $(\lambda_{j})_{j\in\mathbb{Z}}\subset\mathbb{C}\setminus\{0\}$. The following conditions are equivalent:
  1. (i)
    $S^{\ast}$ is $m$-isometric,
  2. (ii)
    $S$ is unitary.
Proof.
It suffices to prove that (i) implies (ii). We have $\mathbf{B}(\mathbb{C})=\mathbb{C}$ (up to an isomorphism). Because the only invertible elements of $\mathbb{C}[z]$ are constant polynomials, it follows that in Theorem 4.7(ii), it has to be $p=1$, so $S$ is isometric. Since, by (2.8), $S$ is invertible, we get that $S$ is unitary. ∎
In general, the conclusion of Corollary 4.8 is not true if $\dim H\geqslant 2$ as is shown in the following example.
Example 4.9.
Define $p\in\mathbf{B}(\mathbb{C}^{2})_{2}[z]$ by the formula:
$p(z)=\begin{bmatrix}1&z\\ z&z^{2}+1\end{bmatrix},\quad z\in\mathbb{C}.$
It is a matter of routine to verify that $p(n)$ is positive and invertible for every $n\in\mathbb{Z}$. Moreover, for every $n\in\mathbb{Z}$ the characteristic polynomial of $p(n+1)p(n)^{-1}$ is equal to $h(z)=z^{2}-3z+1$. Hence,
$r(p(n+1)p(n)^{-1})=r(p(1)p(0)^{-1}),\quad n\in\mathbb{Z}.$
Therefore, $\sup\left\{r(p(n+1)p(n)^{-1})\colon n\in\mathbb{Z}\right\}<\infty$. By Remark 4.5 and Theorem 4.4 there exists a $3$-isometric bilateral weighted shift $S\in\mathbf{B}(\ell^{2}(\mathbb{Z},\mathbb{C}^{2}))$ with positive and invertible weights $(S_{j})_{j\in\mathbb{Z}}$ satisfying (4.1) and (4.2). Set
$q(z):=\begin{bmatrix}z^{2}+1&-z\\ -z&1\end{bmatrix},\quad z\in\mathbb{C}.$
Then $q\in\mathbf{B}(\mathbb{C}^{2})_{2}[z]$ and $p\cdot q=q\cdot p=I$. Hence, by Theorem 4.7, $S^{\ast}$ is also 3-isometric.